The average amount of milligrams (mg) of sodium in the water in a certain neighbourhood in Winnipeg is 60 mg/l, and the standard deviation is 10 mg/l. Assume the variable is normally distributed.
(a) If a sample of water is selected form the neighbourhood, what is the probability that the sodium content will be more than 70 mg/l?
(b) If 10 samples of the water from the neighbourhood is selected, what is the probability that the mean of the sample will be larger than 70 mg/l?
(c) Why is the probability for part (a) greater than that for part (b)?
1
Expert's answer
2020-03-13T11:02:53-0400

"\mu= 60, \sigma=10"

(a) P(X>70)

X is normally distributed with mean 60 and standard deviation 10.

"P(X>70)=1-P(X\le70)"

"Z-value =\frac{X-\mu}{\sigma}"

"=\frac{70-60}{10}=1". From Z table or =NORM.DIST(70,60,10,TRUE),

"P(Z\le1)= 0.841345"

Thus, "P(X>70)=1-0.841345=0.158655"

(b). "P(\bar{X}>70)"

"\bar{X}" Is normally distributed with mean 60 and variance "\frac{10}{\sqrt10}"

"P(\bar{X}>70)=1-P(\bar{X}\le70)"

"Z-value =\frac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}"

"=\frac{70-60}{\frac{10}{\sqrt{10}}}=3.162278"

From Z table or =NORM.DIST(70,60,10/10^0.5,TRUE),"P(\bar{X}\le70)=" 0.999217

Thus, "P(\bar{X}>70)=1-0.999217=0.000783"

(c). As the sample size increases, the sample mean tends to be closer to the population mean, and standard deviation decreases as the sample size increases. Part (a) has a sample size of 1 and the standard deviation is 10, while part (b) has a sample size of 10 leading to the standard deviation of"\frac{10}{\sqrt{10}}" which is smaller than the standard deviation in part (a). Therefore, part (a) has a greater probability since the distribution around X is wider than the distribution around "\bar{X}" in part (b).